# Novel Direct Alpha Spectroscopy Technique for 225Ac Radiopharmaceutical detection in Cancer Cells

### v1

@@ Abstract
Targeted alpha-particle therapy (TAT) employs alpha-emitting radionuclides conjugated to tumor-targeting molecules to deliver localized radiation to cancer cells, showing great promise in treating metastatic cancers. Among these radionuclides, Actinium-225 (225Ac, t1/2 = 9.9 days) has emerged as a clinically promising candidate. Its decay chain generates four successive alpha emissions, resulting in highly localized and effective cytotoxic damage to cancer cells when delivered to tumor sites. However, the assumption of complete retention of 225Ac and its radioactive daughters at these target sites is often inaccurate. The nuclear recoil effect can lead to off-target distribution and unintended toxicity.
To address this challenge, we developed a novel direct alpha spectroscopy method utilizing the Bio-Sample Alpha Detector (BAD). Alpha emission spectra were acquired under atmospheric pressure at room temperature, enabling direct observation of 225Ac and its daughters’ behavior in radiopharmaceutical applications. AR42J rat pancreatic tumor cells, expressing somatostatin receptor 2 (SSTR2), were incubated with [225Ac]Ac-crown-TATE, [225Ac]Ac-PSMA-617, and [225Ac]Ac3+. The BAD setup allowed placement of radiolabeled cell samples within 100 $\mu$m of the detector, facilitating alpha spectra acquisition with statistical uncertainties $<$1%.
Our results revealed distinct spectral differences between radiolabeled cells and reference samples, demonstrating [225Ac]Ac-crown-TATE uptake by AR42J cells. Detection of 213Po, one of the 225Ac decay daughters, highlighted partial retention and release of decay products from cells, providing information on intracellular retention and daughter redistribution. Geant4 simulations confirmed the alignment of experimental data with theoretical models, validating the method’s accuracy. This study establishes a direct alpha spectroscopy approach for investigating 225Ac and its daughters’ behavior in cells and offers a powerful tool for microdosimetry estimation.
@@ 1 Introduction
The field of radiopharmaceuticals has witnessed significant advancements in recent years, particularly in the development of alpha-emitting isotopes for Targeted Alpha Therapy (TAT). TAT leverages the high linear energy transfer (LET) and short range of alpha particles to selectively destroy cancer cells while minimizing damage to surrounding healthy tissue [1, 2, 3].This approach shows particular promise for treating metastatic and micro-metastatic cancers, which often prove resistant to conventional therapies [4]. Alpha particles, characterized by energies typically ranging from 3 to 10 MeV, exhibit a high LET and a short range in biological tissues ($<$100 $\mu$m). These properties enable the delivery of highly localized and lethal radiation doses, inducing severe DNA damage in targeted cells. However, these same properties pose challenges in research, as their short range complicates direct detection, particularly in biological environments where attenuation and scattering can hinder accurate measurements [5]. One alpha-emitting isotope that has gained considerable attention for its therapeutic properties is Actinium-225 (225Ac). 225Ac has shown significant promise for clinical applications due to its relatively long half-life of approximately 10 days. This provides a potentially suitable balance between allowing sufficient time for targeted delivery and sustaining therapeutic activity. The decay chain of 225Ac includes the emission of four alpha particles with energies ranging from 5.8 MeV to 8.4 MeV, each of which is highly effective in inducing double-strand breaks in the DNA of cancer cells, ultimately leading to cell death. This makes 225Ac-labeled radiopharmaceuticals particularly effective in treating cancers such as prostate cancer and neuroendocrine tumors [2, 6]. Despite its advantages, the application of 225Ac-based therapies faces key challenges, particularly in the detection and measurement of alpha emissions in biological systems [6]. One significant challenge is that 225Ac-labeled complexes contain up to seven radioisotopes in their decay chain. The high recoil energy of alpha decay can break coordination bonds, leading to the release of uncomplexed daughter radionuclides. These progenies may redistribute in vivo, making accurate microdosimetry estimation challenging . The redistribution of these daughter radionuclides is still not well understood, adding complexity to dosimetry and potential toxicity assessments [7].
@@
Conventional measurement techniques often rely on indirect gamma detection [8], which is problematic for isotopes like 225Ac that emit alpha particles without easily detectable gamma rays. This limitation prevents real-time monitoring of the behavior and function of alpha emitters in biological systems. Furthermore, conventional alpha detectors, while effective in industrial and scientific settings, are not well-suited for use in biological studies. These detectors typically operate under a vacuum to minimize attenuation, making them incompatible with measurements in living cell cultures. Addressing these challenges requires innovative detection methods that can provide direct and real-time insights into alpha-emitting isotopes under biologically relevant conditions.
@@
This study aims to develop a Bio-Sample Alpha Detector (BAD) and novel measurement protocols that enable in vitro investigation of 225Ac-labeled pharmaceuticals. As a spin-off of a research project related to medical isotopes, we leveraged expertise and technology from the characterization of radioactive ion beams to develop the BAD for direct alpha spectroscopy on cell cultures. By focusing on the development of a silicon PIN (SiPIN) sensor-based alpha detector capable of operating at ambient temperature and normal atmospheric conditions, we investigated the alpha spectra of pancreatic cancerous cells incubated with labeled 225Ac. This approach represents a significant step forward in the field, as it not only enhances our understanding of the behavior of alpha-emitting radiopharmaceuticals in a biological context but also reduces the reliance on animal models. Such advancements contribute to a more ethical and potentially more accurate assessment of therapeutic efficacy.
@@ 2.2 Bio-sample Alpha Detector
The BAD utilizes a HAMAMATSU S-series Si PIN photodiode with an active area of 18 × 18 mm [12]. This photodiode was connected to a Mesytec MSI-8 preamp-shaping amplifier module. A stable bias voltage of +70 V was supplied to the detector using a high voltage power supply from Mesytec. The processed signal from the photodiode was then sent to a 12-bit multi-channel analyzer (Toivel ADC/MCA). Data acquisition was managed with a custom software package using CERN’s ROOT data analysis framework and MIDAS that was adapted for this specific application [13]. For optimal performance, the design of the sample cups was critical. We used SpectroMicro XRF sample cups with dimensions of 23.9 mm × 18.4 mm (outer diameter) × 19.4 mm (height), equipped with a 2.5 µm thick Mylar foil layer ( 1(a)). The choice of Mylar foil was crucial, as its thinness allows alpha particles to pass through with minimal energy loss, ensuring accurate detection. Additionally, the detector setup was specifically designed to minimize the distance between the radioactive source and the detector, achieving a source-to-detector separation of approximately 100 µm. The detection system was then calibrated for energy and efficiency using a 225Ac calibration source prepared by drying a 225Ac solution of known activity in SpectroMicro XRF sample cups. A CAD schematic view of the detector assembly is shown in  1(b).
Figure 1: (a): SpectroMicro XRF sample cup with a Mylar foil layer at the bottom, used for cell sample preparation in direct alpha spectroscopy. The cup dimensions are 23.9 mm $\times$ 18.4 mm (outer diameter) $\times$ 19.4 mm (height), with a Mylar foil thickness of 2.5 $\mu$m to ensure optimal alpha particle detection. (b): CAD 3D cross section view of the detector assembly.
@@
A mass-separated RIB of 225Ac/225Ra isotopes is collected at the ISAC Implantation Station, as detailed in [14]. The 225Ac is then used either directly or extracted from a 225Ra generator for experiments in this study [15]. In addition to ISAC, 225Ac was also sourced from the TRIUMF 520 MeV Isotope Production Facility (IPF). Here, 225Ac was produced by irradiating 232Th foils with a $\sim$ 480 MeV proton beam, with a total irradiation dose of up to 12,500 µAh [16, 10].
@@ 2.4 Radiolabeling of 225Ac-Labeled Crown-TATE and PSMA-617
The molar activity of the radiolabeled compounds was determined using the ratio of the activity of 225Ac to the number of moles of targeting agent used for labeling. In this study, a molar activity of 256 kBq/nmol was achieved for both crown-TATE and PSMA-617 and used for subsequent experiments.
225Ac-labeled compounds were diluted in PBS solution to reach the desired volume for the experiment. The diluted activities were further quantified via gamma spectroscopy using a high-purity germanium (HPGe) detector.
@@ 2.5.1 Culturing and Maintenance of AR42J Cells
The AR42J cell line is an epithelial-like cell isolated from a rat’s pancreas with a tumor [19]. The cells were cultured in a mixture of 60% (v/v) Ham’s F-12K (Kaighn’s) Medium and 40% (v/v) Dulbecco’s Modified Eagle’s Medium (DMEM) on 60 mm plates. The cultures were incubated at 37 °C in a 5% CO2 environment. Both media were supplemented with 10% (v/v) Fetal Bovine Serum (FBS) and 1% (v/v) Penicillin-Streptomycin (PS) prior to being combined in a 60/40 ratio. AR42J cells are adherent, attaching firmly to the bottom of the culture plates, and exhibit a high density of somatostatin receptor 2 (SSTR2) on their membranes, allowing them to specifically recognize and bind to the crown-TATE peptide. The culture medium was refreshed every two days, and cells were passaged weekly upon reaching approximately 80% confluency. For passaging, the spent medium was first removed, and the cells were washed with 10 mL of PBS to remove residual serum. The cells were then detached by adding 2 mL of trypsin and incubating at 37 °C for 5 minutes. Trypsinization was neutralized by adding 8 mL of media, and the resulting cell suspension was transferred to a Falcon tube. The tube was centrifuged at 1300 RPM for 6 minutes. Following centrifugation, the supernatant containing the media and trypsin was discarded, and the cell pellet was resuspended in 10 mL of fresh media. Finally, the cells were seeded onto 100 mm plates with an appropriate dilution factor, typically a 1:2 ratio if passaged every seven days, depending on the confluency of the original plate.
@@ 2.5.2 Preparation of Cell Samples for Direct Alpha Spectroscopy
Figure 2 shows Workflow for the preparation and measurement of AR42J cell samples for alpha spectroscopy. The process began with the removal of the cell culture dish containing the AR42J cell line from the CO2 incubator. The supernatant media was removed, and the cells were washed with 10 mL of PBS. Then, the cells were detached from the plate by adding 2 mL of trypsin and incubating at 37 °C for 5 minutes. The trypsin was then neutralized using 8 mL of media, the cell suspension was transferred to a Falcon tube, and the Falcon tube was taken to the centrifuge. The cells were centrifuged for 6 minutes at 1300 RPM. After centrifugation, the media/trypsin supernatant solution was removed, and the cells were resuspended in 5 mL of fresh media. An aliquot of the cell suspension was taken, and the cells were counted using the Countess FL 3. Based on the cell concentration obtained, 300,000 cells were seeded in each well of a 24-well plate. The cells were incubated at 37 °C for two days to allow sufficient time for cell adhesion and monolayer formation prior to further experimental procedures.
@@
On the day of the experiment, the media (F12K and DMEM) in each well was carefully removed without disturbing the adherent cell monolayer. 400 $\mu$L of the labeled activity, which was diluted in PBS, was added to each well, and the cells were incubated at 37 °C for an hour. Following incubation, the supernatant containing unbound radiolabel was removed, and 200 $\mu$L of trypsin was added to each well to detach the cells from the surface. The well plate was kept in the incubator for 5 minutes to facilitate cell detachment. To neutralize the trypsin, 800 $\mu$L of fresh media was added to each well, bringing the total volume to 1 mL. The cell suspension was then transferred to an Eppendorf tube and centrifuged at 10 G for 6 minutes. After centrifugation, the supernatant was collected in another Eppendorf tube for further analysis, and 400 $\mu$L of fresh media was added to resuspend the cell pellet. The final cell suspension was thoroughly mixed and transferred to SpectroMicro XRF sample cups, which had been pre-coated with poly-D-lysine to promote cell adhesion. It is worth to mentioning Simpler methods such as growing cells on Mylar foil directly, proved inefficient due to 225Ac’s adherence to Mylar, motivating the development of this refined protocol.
@@
Figure 2: Workflow for the preparation and measurement of AR42J cell samples for alpha spectroscopy.
@@
2.5.3 Direct Alpha Spectroscopy of Cell Cultures Incubated with Labeled Activity

### v2

@@ Abstract
Objective: Targeted alpha-particle therapy (TAT) is a promising approach for treating metastatic cancers, utilizing alpha-emitting radionuclides conjugated to tumor-targeting molecules. Actinium-225 (225Ac) has emerged as a clinically relevant candidate due to its decay chain, which produces four successive alpha emissions, effectively damaging cancer cells. However, the nuclear recoil effect can lead to off-target redistribution of decay daughters, complicating dosimetry and increasing potential toxicity. This study aims to address these challenges by developing a direct alpha spectroscopy method for in vitro investigations of 225Ac radiopharmaceuticals. Approach: We developed the Bio-Sample Alpha Detector (BAD), a silicon-based detector designed to operate under ambient conditions, enabling direct alpha spectroscopy of cell samples. AR42J rat pancreatic tumor cells, which express somatostatin receptor 2 (SSTR2), were incubated with [225Ac]Ac-crown-TATE, [225Ac]Ac-PSMA-617, and [225Ac]Ac3+. The BAD setup allowed radiolabeled cell samples to be positioned within 100 $\mu$m of the detector for alpha spectra acquisition with statistical uncertainties of less than 1% in count rates. Geant4 Monte Carlo simulations were employed to validate the experimental results. Main Results: Distinct spectral differences between radiolabeled cells and reference samples confirmed the uptake of [225Ac]Ac-crown-TATE by AR42J cells. Detection of 213Po, a decay daughter of 225Ac, indicated partial retention and release of decay products from cells, providing insight into intracellular retention and radionuclide redistribution. Geant4 simulations confirmed the alignment of experimental data with theoretical predictions. Significance: This study introduces a novel method for directly measuring the behavior of 225Ac and its decay daughters in biological samples using alpha spectroscopy. The BAD setup provides a valuable tool for investigating radionuclide retention, redistribution, and microdosimetry in radiopharmaceutical research.
@@ 1 Introduction
The field of radiopharmaceuticals has witnessed significant advancements in recent years, particularly in the development of alpha-emitting isotopes for Targeted Alpha Therapy (TAT). TAT leverages the high linear energy transfer (LET) and short range of alpha particles to selectively destroy cancer cells while minimizing damage to surrounding healthy tissue [1, 2, 3].This approach shows particular promise for treating metastatic and micro-metastatic cancers, which often prove resistant to conventional therapies [4]. Alpha particles, characterized by energies typically ranging from 3 to 10 MeV, exhibit a high LET and a short range in biological tissues ($<$100 $\mu$m). These properties enable the delivery of highly localized and lethal radiation doses, inducing severe DNA damage in targeted cells. However, these same properties pose challenges in research, as their short range complicates direct detection, particularly in biological environments where attenuation and scattering can hinder accurate measurements [5]. One alpha-emitting isotope that has gained considerable attention for its therapeutic properties is Actinium-225 (225Ac). 225Ac has shown significant promise for clinical applications due to its relatively long half-life of approximately 10 days. This provides a potentially suitable balance between allowing sufficient time for targeted delivery and sustaining therapeutic activity. The decay chain of 225Ac includes the net emission of four alpha particles with energies ranging from 5.8 MeV to 8.4 MeV, each of which is highly effective in inducing double-strand breaks in the DNA of cancer cells, ultimately leading to cell death. This makes 225Ac-labeled radiopharmaceuticals particularly effective in treating cancers such as prostate cancer and neuroendocrine tumors [2, 6].
Despite its advantages, the application of 225Ac-based therapies faces key challenges, particularly in the detection and measurement of alpha emissions in biological systems [6]. One significant issue is that 225Ac-labeled complexes contain up to seven radioisotopes in their decay chain. The high recoil energy from alpha decay can disrupt chemical bonds, releasing daughter radionuclides that may redistribute in vivo. Although several in vivo studies have mapped the biodistribution of key daughter isotopes such as Francium-221 (221Fr), Astatine-217 (217At), and Bismuth-213 (213Bi), these approaches do not provide direct alpha particle measurements. The redistribution patterns have been shown to vary depending on the pharmacokinetics of the targeting vector, with slower-clearing antibodies releasing daughters into circulation and faster-clearing small molecules directing them toward the liver or kidneys [7, 8]. These findings have significantly improved dosimetry models and highlighted the importance of accounting for daughter migration in absorbed dose estimates. However, most of this knowledge derives from in vivo models, and there remains a need for complementary in vitro methods that allow for direct, real-time assessment of radiopharmaceutical behavior and retention at the cellular level.
@@
Conventional measurement techniques often rely on indirect gamma detection [9], which is problematic for isotopes like 225Ac that emit alpha particles without easily detectable gamma rays. This limitation prevents real-time monitoring of the behavior of alpha emitters in biological systems. Furthermore, conventional alpha detectors, while effective in industrial and scientific settings, are not well-suited for use in biological studies. These detectors typically operate under a vacuum to minimize attenuation, making them incompatible with measurements in living cell cultures. Addressing these challenges requires innovative detection methods that can provide direct and real-time insights into alpha-emitting isotopes under biologically relevant conditions.
@@
This study aims to develop a Bio-Sample Alpha Detector (BAD) and novel measurement protocols for in vitro investigation of 225Ac-labeled pharmaceuticals. As a spin-off of a research project related to medical isotopes, we leveraged expertise and technology from the characterization of radioactive ion beams to develop the BAD for direct alpha spectroscopy on cell cultures. By focusing on the development of a silicon PIN (Si-PIN) sensor-based alpha detector capable of operating at ambient temperature and normal atmospheric conditions, we investigated the alpha spectra of pancreatic cancerous cells incubated with labeled 225Ac. While this approach provides valuable insights into the behavior of alpha-emitting radiopharmaceuticals at the cellular level, a full understanding of the redistribution and biological impact of recoiling daughter radionuclides will require further evaluations in more complex biological systems, such as animal models. Nonetheless, this study represents an important first step in advancing detection methods for TAT and reducing reliance on animal studies for initial evaluations of radionuclide behavior.
@@ 2.2 Bio-sample Alpha Detector
The BAD utilizes a HAMAMATSU S-series Si-PIN photodiode with an active area of 18 × 18 mm [13]. This photodiode was connected to a Mesytec MSI-8 preamp-shaping amplifier module. A stable bias voltage of +70 V was supplied to the detector using a high voltage power supply from Mesytec. The processed signal from the photodiode was then sent to a 12-bit multi-channel analyzer (Toivel ADC/MCA). Data acquisition was managed with a custom software package using CERN’s ROOT data analysis framework and MIDAS that was adapted for this specific application [14]. For optimal performance, the design of the sample cups was critical. We used SpectroMicro XRF sample cups with dimensions of 18.4 mm (outer diameter), 12.1 mm (inner diameter), and 19.4 mm (height), equipped with a 2.5 µm thick Mylar foil layer ( Figure 1). The choice of Mylar foil was crucial, as its thinness allows alpha particles to pass through with minimal energy loss, ensuring accurate detection. Additionally, the detector setup was specifically designed to minimize the distance between the radioactive source and the detector, achieving a source-to-detector separation of approximately 100 µm. The detection system was then calibrated for energy and efficiency using a 225Ac calibration source prepared by drying a 225Ac solution of known activity in SpectroMicro XRF sample cups. A Computer-Aided Design (CAD) schematic view of the detector assembly is shown in  Figure 1. The activity in the 225Ac calibration source was 4 kBq at the time of measurement. The reported activity accounts for both 225Ac and its decay daughters, as the measurement was performed after secular equilibrium had been reached.
[…]
Figure 1: (a): SpectroMicro XRF sample cup with a Mylar foil layer at the bottom, used for cell sample preparation in direct alpha spectroscopy. The cup inner diameter is 12.1 mm, with a Mylar foil thickness of 2.5 $\mu$m to ensure optimal alpha particle detection. (b): CAD 3D cross section view of the detector assembly.
@@
A mass-separated RIB of 225Ac/225Ra isotopes is collected at the ISAC Implantation Station, as detailed in [15]. The use of electromagnetic mass separation ensures the isolation of 225Ac with high isotopic purity, effectively excluding 227Ac contamination. The 225Ac is then used either directly or extracted from a 225Ra generator for experiments in this study [16]. In addition to ISAC, 225Ac was also sourced from the TRIUMF 520 MeV Isotope Production Facility (IPF). Here, 225Ac was produced by irradiating 232Th foils with a $\sim$ 480 MeV proton beam, with a total irradiation dose of up to 12,500 µAh [17, 11], However, the 225Ac used in this study was obtained from a second-pass separation via a 225Ra generator, which yields 225Ac with negligible or no 227Ac contamination, as demonstrated in [11].
@@ 2.4 Radiolabeling of 225Ac-labeled crown-TATE and PSMA-617
The molar activity of the radiolabeled compounds was determined using the ratio of the activity of 225Ac to the number of moles of targeting agent used for radiolabeling. In this study, a molar activity of 256 kBq/nmol was achieved for both crown-TATE and PSMA-617 and used for subsequent experiments. This corresponds to a specific activity of approximately 0.168 kBq/µg for crown-TATE (molecular weight: 1522 g/mol) and 0.202 kBq/µg for PSMA-617 (molecular weight: 1267 g/mol). The ligand-to-metal ratio of 1903:1 indicates that, on average, one 225Ac atom was coordinated by approximately 1900 targeting peptides in both cases during the radiolabeling process.
225Ac-labeled compounds were diluted in DPBS solution to reach the desired volume for the experiment. The diluted activities were further quantified via gamma spectroscopy using a high-purity germanium (HPGe) detector. The activity levels used for each sample type ranged from 3 to 10 kBq. A summary of the radiolabeling conditions and properties for the 225Ac-labeled compounds is presented in Table 1.
Table 1: Radiolabeling conditions for each compound used in cell incubation studies.
Compound Temp. (°C) Time (min) RCY (%) Molar Activity (kBq/nmol)
[…]
@@ 2.5.1 Culturing and maintenance of AR42J cells
The AR42J cell line, derived from a rat pancreatic tumor [20], was cultured using standard adherent cell culture techniques. Cells were maintained in a 60:40 mixture of F-12K and DMEM media, supplemented with 10% FBS and 1% PS, at 37 °C in a humidified incubator with 5% CO2. Cells were seeded onto 60 mm tissue culture plates and passaged weekly upon reaching approximately 80% confluency. The number of passages was kept below 20. For passaging, the medium was aspirated, and the cells were washed with DPBS to remove residual serum. Cells were then detached by adding 2 mL of trypsin-EDTA solution and incubated for 5 minutes at 37 °C. Trypsin was neutralized with 8 mL of complete media, and the cell suspension was transferred to a Falcon tube for centrifugation at 1300 RPM for 6 minutes. The resulting pellet was resuspended in fresh media and reseeded at a 1:2 dilution, depending on confluency. AR42J cells are strongly adherent and express a high density of somatostatin receptor 2 (SSTR2), allowing for specific binding of crown-TATE targeting peptides.
@@ 2.5.2 Preparation of cell samples for direct alpha spectroscopy
Table 2 shows the Workflow for the preparation and measurement of AR42J cell samples for alpha spectroscopy. The supernatant media was removed, and the cells were washed with 10 mL of DPBS. Then, the cells were detached from the plate by adding 2 mL of trypsin and incubating at 37 °C for 5 minutes. The trypsin was then neutralized using 8 mL of media, the cell suspension was transferred to a Falcon tube, and the Falcon tube was taken to the centrifuge. The cells were centrifuged at 1300 RPM (corresponding to approximately 317 g) for 6 minutes. After centrifugation, the media/trypsin supernatant solution was removed, and the cells were resuspended in 5 mL of fresh media. An aliquot of the cell suspension was taken, and the cells were counted using the Countess FL 3. Based on the cell concentration obtained, 300,000 cells were seeded in each well of a 24-well plate. The cells were incubated at 37 °C for two days to allow sufficient time for cell adhesion and monolayer formation prior to further experimental procedures. After two days, the number of cells was found to be $4.19\times 10^{5}$, with a viability of 85%.
@@
On the day of the experiment, the media (F12K and DMEM) in each well was carefully removed without disturbing the adherent cell monolayer. 400 $\mu$L of the labeled activity, which was diluted in DPBS, was added to each well, and the cells were incubated at 37 °C for one hour. Following incubation, the supernatant containing unbound radiolabel was removed, and 200 $\mu$L of trypsin was added to each well to detach the cells from the surface. The well plate was kept in the incubator for 5 minutes to facilitate cell detachment. To neutralize the trypsin, 800 $\mu$L of fresh media was added to each well, bringing the total volume to 1 mL. The cell suspension was then transferred to an Eppendorf tube and centrifuged at 10 G for 6 minutes. After centrifugation, the supernatant was collected in another Eppendorf tube for further analysis, and 400 $\mu$L of fresh media was added to resuspend the cell pellet. The final cell suspension was thoroughly mixed and transferred to SpectroMicro XRF sample cups, which had been pre-coated with poly-D-lysine to promote cell adhesion. It is worth mentioning that simpler methods, such as growing cells directly on Mylar foil, proved inefficient due to 225Ac’s adherence to Mylar foil, which motivated the development of this refined protocol.
@@
Table 2: Workflow for the preparation and measurement of AR42J cell samples for alpha spectroscopy.
Step Procedure
Two days prior to experiment Prepare sample cups with Mylar foil and coat with poly-D-lysine Seed cells to 24-well plate Incubate at 37 °C for two days
Cell sample preparation Remove and discard media above the cell monolayer Add 400 µL of labeled activity diluted in DPBS Incubate cells at 37 °C for 1 hour Remove and discard supernatant Add 200 µL of trypsin Incubate for 5 minutes to facilitate cell detachment Add 800 µL of fresh media to neutralize trypsin and bring volume to 1 mL Transfer suspension to Eppendorf tube and centrifuge at 10 G for 6 minutes Collect supernatant for further measurement and analysis Add 400 µL of fresh media to pellet and mix thoroughly Transfer final suspension to sample cup
Alpha spectrum measurement Measure cell sample for 900 seconds
@@
2.5.3 Direct alpha spectroscopy of sell cultures incubated with labeled activity